Please help me solve this equation by using substitution, Bob is one third of the age of his father. In 12 years he will be half the age of his father. How old is each now? Thank you.
Answer by scott8148(6628)

"Bob is one third of the age of his father" ___ B = (1/3)f

"In 12 years he will be half the age of his father" ___ B + 12 = (1/2)(f + 12)

substitute the equivalent value of B from the 1st equation into the 2nd equation and find "f"

then substitute the value of "f" into the first equation to find B